OpenStudy (magbak):

PLEASE PLEASE PLEASE PLEASE PLEASE PLEASE HELP ME!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!! Medal will be awarded. Using complete sentences, explain how to graph the following piecewise function: y = x2 if x < 1 y = 3x if x greater than or equal to 1

OpenStudy (magbak):

Thank you for your time.

jimthompson5910 (jim_thompson5910):

y = x^2 if x < 1 so you graph y = x^2, but you only graph it for values of x less than 1 (so you erase the portion to the right of x = 1)

jimthompson5910 (jim_thompson5910):

similarly, you plot y = 3x, but only for x values that are greater than or equal to 1

OpenStudy (magbak):

That is it

jimthompson5910 (jim_thompson5910):

yeah basically

OpenStudy (magbak):

OK thank you. But one thing is their a more extraneous explanation.

jimthompson5910 (jim_thompson5910):

what do you mean

OpenStudy (magbak):

A longer more detailed solution or is that smack up just it.

jimthompson5910 (jim_thompson5910):

you could go in more detail, but that's the basics of it

jimthompson5910 (jim_thompson5910):

you could explain that x^2 is plotted from a series of points and so is y = 3x

jimthompson5910 (jim_thompson5910):

y = 3x only needs 2 points but x^2 needs about 5 or 6 to get an idea of the shape of the graph

OpenStudy (magbak):

If it is not to much I would prefer the longer explanation.

jimthompson5910 (jim_thompson5910):

you could do it like this: use these x values and plug them into the piece wise function -3, -2, -1, 0, 1, 2, 3 the values -3, -2, -1, 0 will go into y = x^2 to get corresponding y values so you'll have a bunch of points which will then be connected with a curve to graph y = x^2

jimthompson5910 (jim_thompson5910):

the values x = 1, x = 2, x = 3 will go into y = 3x and they will generate points to help graph y = 3x

OpenStudy (magbak):

Thank you.

OpenStudy (magbak):

Ok that is perfect thank you.

OpenStudy (magbak):

Was this a good question.

jimthompson5910 (jim_thompson5910):

what do you mean?

OpenStudy (magbak):

Was it a rare type. You do not find piece wise functions a lot do you.

jimthompson5910 (jim_thompson5910):

oh I've seen a lot of piece wise function questions, so not really

jimthompson5910 (jim_thompson5910):

but today is the first time I've seen it, so I guess it's not as common

OpenStudy (magbak):

Oh well I thought it was really rare.?!

jimthompson5910 (jim_thompson5910):

hmm not really

OpenStudy (magbak):

Thank you. One thing if it is not a lot can I get a medal for asking please.

jimthompson5910 (jim_thompson5910):

sure

OpenStudy (magbak):

Thank you

OpenStudy (magbak):

Bye